OpenStudy (anonymous):

Find the global maximum and minimum for the function on the closed interval. f(x)=x^3 - 3x^2 + 2 1≤x≤3 Write the x values in increasing order. The global minimum at x=_____ and x=____ is _____. The global maximum at x=_____ and x=____ is _____. please explain! thank you:)

ganeshie8 (ganeshie8):

find all the local max/mins first

OpenStudy (anonymous):

i do not by finding the derivatives first right? f ' (x)= 3x^2 - 6x and f '' (x)= 6x-6 ?

ganeshie8 (ganeshie8):

yes, the slope is 0 at max/min points

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

so set the first derivative to 0 and solve x

ganeshie8 (ganeshie8):

f ' (x)= 3x^2 - 6x = 0 x = ?

OpenStudy (anonymous):

okay so 3x(x-6)=0 x=0; x=6 ?

ganeshie8 (ganeshie8):

yes, 0 and 6 are the critical points

ganeshie8 (ganeshie8):

u familiar wid second derivative test ?

ganeshie8 (ganeshie8):

it is not needed for doing this problem, but if you're familiar wid that, we can use it...

OpenStudy (anonymous):

is it where you use the signs?  like this?